# Johannes Huemer [^2] These authors contributed equally to this work. Markus Murschitz33footnotemark: 3 Matthias Schörghuber Lukas Reisinger Thomas Kadiofsky Christoph Weidinger Mario Niedermeyer Benedikt Widy Marcel Zeilinger Csaba Beleznai Tobias Glück Andreas Kugi Patrik Zips44footnotemark: 4

### v1

@@ 5.3.3 Manipulation
The Manipulation action involves the correct pick-up and drop-off of the pallets. The high-level task planning framework controls this action by selecting a suitable target pallet based on spatial proximity and accessibility and then executing a controlled approach. To ensure correct pallet pickup, both the vehicle base and fork cascaded control loops from Section 5.3.1 are necessary. The aim is to minimize the error between the pallet pose and the vehicle’s rear axle. The target pose for the vehicle base control is now the center of the front face of the pallet ($\mathbf{q}^{d}=\mathbf{q}_{P}$) and the vehicle origin is the center of the rear axle $\mathbf{q}_{R}$ [^6] Remember that the forks are mounted at the rear of the vehicle., as illustrated in Figure 19(b).
During the pallet approach phase, the FTT component from Figure 16b is crucial to align the fork tips with the pallet openings, before and while the vehicle is moving onto the pallet. The reverse action of placing the pallet follows a process similar to pickup. Here, a virtual pallet called slot defines the target location, guiding the system to the desired position. Figure 20 illustrates data from a successful pallet manipulation maneuver, showing the process of placing a pallet into a slot on the truck. The critical moment for loading the pallet, highlighted by a dashed gray line, occurs when the fork tips reach the front of the (virtual) pallet, and the forks must properly engage with the pallet pockets. The vehicle stops moving ($v^{d}=0$) at $x_{P}^{\prime}=-d_{F}$, where $d_{F}$ corresponds to the length of the forks or if the fork-mounted laser sensor indicates successful pallet insertion.
Figure 20: Exemplary data for placing a pallet in a slot on the truck. Cyan lines indicate setpoint signals, red lines measurements, and the gray dashed line marks the time when the fork tip reaches the front of the slot. The standstill time ($\dot{x}^{\prime}_{P}=0$) of $\approx 10$ seconds results from the waiting time required for the forks to reach the desired height.
Initial iterations of ADAPT have shown that placing the pallet on its designated slot purely by estimating its z-position based on the ground plane or a truck loading edge is not sufficiently robust. Thus, we introduce a simple improvement procedure based on measuring the hydraulic pressure in the fork lifting cylinder, as follows:
@@
2:     $crit\_cnt\leftarrow crit\_cnt+1$, $f_{pc}\leftarrow$$true$
@@
4:     $crit\_cnt\leftarrow 0$, $f_{pc}\leftarrow$$false$
@@
6:if $crit\_cnt>crit\_tout$ or $p_{l}<p_{c}$ then
@@
A graphical description based on exemplary pallet unloading measurements is shown in Figure 21. In addition to ensuring that a pallet is placed at the correct height, this method also serves as a safety feature to prevent dangerous downward fork movement in case the forks come into contact with the environment. The threshold levels indicating possible ($p_{pc}$) and definite fork contact ($p_{c}$) are empirically selected as a trade-off between operational robustness and safety.
@@ 6 Validation and Evaluation
This section describes the experimental setup in our outdoor large-scale robotics lab followed by a detailed analysis of the robust long-term operation together with a direct comparison with an experienced human operator. At the end, lessons learnt of the development and testing process are presented to improve development for future work.
@@ 6.1 Experimental Setup
The experiment is designed to evaluate the long-term operational robustness and performance of ADAPT across close-to-real loading scenarios. Three distinct operational modes were tested: ground-to-ground (G to G) pallet loading, ground-to-truck (G to T) pallet loading, and truck-to-ground (T to G) pallet unloading. In this context, the experimental setup involved a truck loaded with four Euro-pallets, each carrying varying types of loads to simulate real-world handling conditions. The initial location for the forklift and the truck parking location are unchanged in all experiments, to allow for a fair comparison between baseline data collected from an expert operator and data collected from autonomous operation. Figure 22 shows the test site with an overlay of the configuration of the ground loading zones (Zone 1 & Zone 2) as well as the truck location.
Figure 22: Satellite image showing the test site with an overlay for the configuration of the loading zones (Zone 1 & Zone 2) as well as the truck location used for the experiments.
@@
As noted above, one of the purposes of this work is to evaluate the automated system under varying weather conditions. Therefore, the baseline data, as well as the data from autonomous operation, are collected under different weather conditions including sunny, cloudy, and rainy weather. All ground-to-ground loading scenarios were carried out in rainy conditions as illustrated in Fig. 24.
@@ 6.3 Performance Criteria
To compare the performance of ADAPT with an experienced operator, manual operation was recorded with the same forklift. The maximum velocity in autonomous operation was limited to the legally maximum allowed velocity of $v_{max}=6$km/h. The following performance criteria are selected for evaluation:
Time Taken for Full Loading Cycle: Total time taken to complete the entire pallet loading and unloading process, from the first pallet pick-up to the last pallet placement. This provides a direct measure of throughput, allowing a clear comparison of speed and productivity between the automated system and the human operator.
Distance Driven: Total distance traveled for the full load task. This enables a comparison between the route efficiency of the automated system’s algorithms and the decision-making of an expert operator. Lower travel distances also equal less energy consumption.
@@ 6.4 Performance Analysis
In the simpler G to G scenario, the autonomous system achieves approximately 88% of expert-level performance, demonstrating its strong potential for timely deployment in commercial operations. In the more complex task of loading onto the truck, the autonomous system remains competitive, showing a time difference of 14% compared to the performance of experts. The largest discrepancy, 26%, is observed in the loading-from-truck scenario. The extended standstill times observed in ADAPT primarily result from computational and detection delays, as well as inefficiencies associated with the sequential execution of tasks, as discussed in detail later.
@@
To evaluate the potential for performance and improvement in more detail, Figure 26 provides a statistical time analysis based on the subtasks of autonomous operation. The missions are defined as discussed in Section 5 as ApproachPallet, LoadPallet, ApproachSlot and UnloadPallet. The two submissions FindPallets and SelectPallets are included in ApproachPallet to increase readability. The two Approach missions consist mostly of typical navigation movement, whereas the load and unload missions deal mostly with manipulation.
@@
Figure 26: Time evaluation of the sub-missions for ground-to-ground (G to G), ground-to-truck (G to T) and truck-to-ground (T to G) loading for autonomous operation (red) and the forklift operator (cyan).
@@
For the ApproachPallet, a longer time and a higher variance for ADAPT can be observed. This arises mostly from two properties of the pallet detection: Firstly, when no pallets are known to the system, ADAPT stops for several seconds in front of the loading zone to obtain a stable object list. Secondly, as the position and orientation of the pallets are not known in advance, a new approach of pallets is sometimes mandatory for correct loading which takes up to 30 seconds. Especially for the G to G scenario, this occurred multiple times resulting in the larger variance for the approach. In the G to T case, the performance of ADAPT and the expert is very similar because the expert operator chooses a suboptimal path as detailed in the latter. For the ApproachSlot mission, ADAPT achieves a very similar performance, even though more changes in driving direction were made.
@@
For the manipulation subtasks, ADAPT performs very consistently for all scenarios. For loading, the performance of G to G and G to T is very similar to that of the human operator. The loading of the truck takes longer for ADAPT. This mostly arises from the time needed to raise the forks to the height of the loading platform and the additional pallet detection check when the pallet is located in a plane different from the movement plane. The unloading of the pallets is close to a standardized task and thus can be done with accurate timing with hardly any variance by the autonomous system. The key difference from manual operation is that the autonomous system raises the forks while stationary, as shown in Figure 20, while a human operator simultaneously manipulates the forks and drives to maximize efficiency. This illustrates a case where ADAPT is not yet optimized for timing performance.
@@
Manual Interventions: During the recorded operation period of more than 3 hours of autonomous functionality, during which 60 pallets were loaded, a total of 17 manual interventions were required. These manual interventions were classified into three categories: Collision Detection, Manipulation and Navigation. The distribution of types of manual intervention is illustrated on the left-hand side of Figure 29.
Figure 29: Overall manual interventions for the autonomous loading operation categorized for their type and their intervention severity.
Notably, nearly 60% of these interventions were attributed to the collision avoidance module, which requires human clearance for any detected potential collision. Many of these collision warnings occurred near the boundary of the map, where the distance from the collision is shorter if approaching an obstacle or boundary of the map. This issue, resulting from stringent safety constraints in the laboratory environment, could potentially be alleviated by implementing obstacle-classification-based warning distances, assigning larger safety margins for humans and moving obstacles, and smaller margins for static obstacles. Additionally, collision warnings were correctly triggered when humans, including the safety operator, were detected in the vehicle’s path, although such incidents were rare. The second category involves inaccuracies in the detection and manipulation modules, which result in failed pallet load and unload attempts. Implementing recovery routines could reduce the number of necessary manual interventions, albeit at the cost of longer execution times. Lastly, navigation and localization interventions occur when no path to a desired position could be planned or when the localization module fails. On the right-hand side of Figure 29, the severity of manual interventions is classified. To this end, three categories of expertise are introduced: 1) Worker for a typical worker on site with little knowledge of the machine, 2) Operator which denotes an educated forklift operator capable of driving the machine manually and 3) Engineer as a robotic engineer with in-depth knowledge of the autonomy framework. Most interventions can be handled by a regular worker by simply ensuring that the corridor in front of the autonomous vehicle is free of obstacles. For ADAPT this is verified by pressing a conformation button once. Approximately one third of interventions require an operator for forklift operations to assist in navigation or manipulation. In our tests, there was one instance where an engineer with knowledge of the autonomous functions was needed, due to a failure in the GNSS localization system. Fig. 30 presents the time analysis of manual interventions during autonomous operation. During a total of 233 minutes of autonomous operation, only 12 minutes were spent in manual control, accounting for less than 5% of the total operation time. Most of this manual intervention, 10 minutes, was dedicated to resolving a GNSS localization issue, which is primarily a prototype-related challenge. All other combined manual interventions accounted for just 2 minutes, representing less than 1% of the total operation time. These results highlight the substantial reduction in manual labor required for material transport, demonstrating the efficiency of ADAPT.
@@
6.5 Lessons Learned
This section outlines key lessons learned from the development and testing of ADAPT, providing insights that we hope will assist fellow researchers in the advancement of autonomous machine technology and the enhancement of the reliability of the system.
All-weather Operation: The initial iteration of our collision avoidance system was based on a wide-angle stereo camera and was tested in an artificial outdoor rain facility designed for autonomous vehicle evaluation. The tests showed that our system performed better in rainy conditions than comparable LiDAR-based solutions, probably due to algorithmic advantages (i.e., a stereo-based rain filter) and mechanical rain protection. For the current iteration, we integrated a LiDAR system with overhead rain protection. Its wider field of view enhances obstacle detection and enables mapping, which stereo cameras alone cannot provide. Additionally, the stereo camera system faced reliability issues in sub-zero temperatures due to USB controller failures, which were mitigated by switching to LiDAR.
In contrast, at high temperatures (above 30°C), we encountered overheating problems with an earlier iteration of the switch cabinet, as its unavoidable placement near the exhaust of the diesel engine exposed it to excessive heat. The current design addresses this problem with an improved inlet for air flow required by the active air-cooling system.
RTK Evaluation: During an earlier system iteration, inconsistent RTK positioning was observed. Extensive investigation revealed that the vehicle’s internal environment contributed significantly to signal interference. Notably, reducing the number of active electronic devices — particularly a USB 3.0 stereo camera connected via a 5-meter cable — mitigated interference. Further research showed that USB 3.0 devices can disrupt GNSS signals, see, e.g., [80]. The replacement of the USB 3.0 cable with a fiber optic cable for data transmission while maintaining a metallic power supply effectively eliminated interference, leading to a significant improvement in the reliability of the RTK positioning.
Multi-sensor Calibration: The factory ZED2i calibration did not meet our needs in terms of accuracy, hence intrinsic camera and stereo calibration have been performed with our proprietary calibration pipeline. In addition, semi-automatic sensor-to-sensor calibrations save time during hardware iterations and can be used to correct aging drift.
Positioning Accuracy: A measurement campaign revealed that in 80% of unloading cases, the global positioning accuracy of ADAPT was sufficient to unload pallets within a 5 cm tolerance without requiring additional local measurements. However, the remaining 20% cases require the continued use of additional local correction mechanisms, such as repeated loading edge detection or local control mechanisms that rely on the LiDAR sensor mounted on a fork.
Simulative Verification and Testing: A significant part of the development and testing of autonomous functionality for ADAPT was carried out in simulation environments. The initial approach aimed for a holistic simulation to test all functionalities in one environment. However, this proved counterproductive, as different tasks have distinct requirements (e.g., photorealism for perception versus real physics for motion control). To address this, specialized simulation tools were used: Gazebo [^7] https://classic.gazebosim.org (accessed 2025-02-28) for task and motion planning and control algorithms, and Blender [^8] https://www.blender.org (accessed 2025-02-28) for object detection evaluation. This approach offered a cost-effective and safe way to validate autonomous functionalities before real-world deployment, especially considering the extensive setup time for prototype tests. Another advantage of simulation was the isolated testing of individual functions, which improved the identification of error sources and was often challenging on the real prototype.
@@ 7 Conclusion and Outlook
This article presented the autonomous outdoor forklift ADAPT, a system capable of operating in the complex and unstructured environments of construction sites, even in challenging conditions, including low to medium rain. By integrating a forklift-specific sensor suite with a tailored software stack, the system achieves fully autonomous pallet loading and transportation. Extensive real-world testing demonstrated that the autonomous forklift operates at near-human efficiency, achieving more than 80% of expert level performance in the demonstrated scenarios, while requiring minimal operator intervention.
A key contribution of this work is the novel factor-graph-based joint optimization approach for vehicle localization and pallet mapping, specifically tailored for object manipulation tasks such as pallet loading. This method enhances the system’s adaptability and accuracy in dynamic construction site environments. Additionally, the integration of a novel fork contact measurement based on pressure feedback significantly improves robustness and safety during object manipulation while remaining cost-effective and simple to implement.
A detailed long-term evaluation compared the autonomous forklift’s performance to that of an expert human operator with over 20 years of experience. The analysis provided insights into performance, robustness, and the nature and frequency of required human interventions. These findings highlight the system’s potential to improve efficiency and safety in material handling for construction sites, addressing labor shortages and reducing operational risks.
Future improvements will focus on motion planning in dynamic environments to enhance safety while maintaining operational efficiency. Additionally, the incorporation of semantic information into the environment mapping approach will complement the existing geometric model, allowing the forklift to prioritize stable surfaces such as concrete roads over more challenging terrains like gravel or soil. Further efforts will be directed toward increasing robustness when navigating through challenging terrain and extreme weather conditions. Recurring hardware updates, particularly those involving processing components, will allow for either replacing existing hardware with more compact, resource-efficient alternatives or utilizing higher-performance components to improve overall timing-related performance. By addressing these aspects, the system can continue to evolve toward a fully integrated, intelligent solution for autonomous material handling in demanding construction environments.

### v2

@@ 5.3.3 Manipulation
The manipulation action involves the correct pick-up and drop-off of the pallets. The high-level task planning framework controls this action by selecting a suitable target pallet based on spatial proximity and accessibility and executing a controlled approach. The vehicle base and fork cascaded control loops from Section 5.3.1 are necessary to ensure correct pallet pick-up. The aim is to minimize the error between the pallet pose and the vehicle’s rear axle. The target pose for the vehicle base control is now the center of the front face of the pallet ($\mathbf{q}^{d}=\mathbf{q}_{P}$), and the vehicle origin is the center of the rear axle $\mathbf{q}_{R}$ [^6] Remember that the forks are mounted at the rear of the vehicle., as illustrated in Figure 20(b).
During the pallet approach phase, the FTT component from Figure 16b is crucial to align the fork tips with the pallet openings, before and while the vehicle is moving onto the pallet. The reverse action of placing the pallet follows a process similar to pick-up. Here, a virtual pallet called slot defines the target location, guiding the system to the desired position. Figure 21 illustrates data from a successful pallet manipulation maneuver, showing the process of placing a pallet into a slot on the truck. The critical moment for loading the pallet, highlighted by a dashed gray line, occurs when the fork tips reach the front of the (virtual) pallet, and the forks must properly engage with the pallet pockets. The vehicle stops moving ($v^{d}=0$) at $x_{P}^{\prime}=-d_{F}$, where $d_{F}$ corresponds to the length of the forks or if the fork-mounted laser sensor indicates successful pallet insertion.
Figure 21: Exemplary data for placing a pallet in a slot on the truck. Cyan lines indicate setpoint signals, red lines refer to measurements, and the gray line marks the time when the fork tip reaches the front of the slot. The standstill time ($\dot{x}^{\prime}_{P}=0$) of $\approx 10$ seconds results from the waiting time required for the forks to reach the desired height.
Initial iterations of ADAPT have shown that placing the pallet on its designated slot purely by estimating its z-position based on the ground plane or a truck loading edge is not sufficiently robust. Thus, we introduce a simple improvement procedure based on measuring the hydraulic pressure in the fork-lifting cylinder, as follows:
@@
2:     $crit_{cnt}\leftarrow crit_{cnt}+1$, $f_{pc}\leftarrow$$true$
@@
4:     $crit_{cnt}\leftarrow 0$, $f_{pc}\leftarrow$$false$
@@
6:if $crit_{cnt}>crit_{tout}$ or $p_{l}<p_{c}$ then
@@
A graphical description based on exemplary pallet unloading measurements is shown in Figure 22. In addition to ensuring that a pallet is placed at the correct height, this method also serves as a safety feature to prevent dangerous downward fork movement in case the forks come into contact with the environment. The parameters indicating possible ($p_{pc}$) and definite fork contact ($p_{c}$,$crit_{tout}$) are empirically selected as a trade-off between operational robustness and safety.
@@ 6 Validation and Evaluation
This section describes the experimental setup in our outdoor large-scale robotics lab followed by a detailed analysis of the robust continuous operation together with a direct comparison with an experienced human operator.
@@ 6.1 Experimental Setup
The experiment is designed to evaluate the operational robustness and performance of ADAPT across close-to-real loading scenarios. Three distinct operational modes were tested: ground-to-ground (G to G) pallet loading, ground-to-truck (G to T) pallet loading, and truck-to-ground (T to G) pallet unloading. In this context, the experimental setup involved a truck loaded with four Euro-pallets, each carrying varying types of loads to simulate real-world handling conditions. The initial location for the forklift and the truck parking location are unchanged in all experiments, to allow for a fair comparison between baseline data collected from an expert operator and data collected from autonomous operation. Figure 23 shows the test site with an overlay of the configuration of the ground loading zones (Zone 1 & Zone 2) as well as the truck location.
Figure 23: Satellite image showing the test site with an overlay for the configuration of the loading zones (Zone 1 & Zone 2) and the truck location used for the experiments.
@@
As noted above, one of the purposes of this work is to evaluate the autonomous system under varying weather conditions. Therefore, the baseline data, as well as the data from autonomous operation, are collected under different weather conditions including sunny, cloudy, and rainy weather. All ground-to-ground loading scenarios were carried out in rainy conditions as illustrated in Fig. 25.
@@ 6.3 Performance Criteria
The manual operation was recorded with the same forklift to compare the performance of ADAPT with that of an expert operator. The maximum velocity in autonomous operation was limited to the legally maximum allowed velocity of $v_{max}=6$km/h. The following performance criteria are selected for evaluation:
Time taken for full loading cycle: Total time taken to complete the entire pallet loading and unloading process, from the first pallet pick-up to the last pallet placement. This provides a direct measure of throughput, allowing a clear comparison of speed and productivity between the autonomous system and the human operator.
Distance driven: Total distance traveled for the full load task. This enables a comparison between the route efficiency of the autonomous system’s algorithms and the decision-making of an expert operator. Lower travel distances also equal less energy consumption.
@@ 6.4 Performance Analysis
In the simpler ground-to- ground scenario, the autonomous system achieves approximately 88% of expert-level performance, demonstrating its strong potential for timely deployment in commercial operations. In the more complex ground-to-truck scenario, the autonomous system remains competitive, showing a time difference of 14% compared to the performance of experts. The largest discrepancy, 26%, is observed in the truck-to-ground scenario. The extended standstill times observed in ADAPT primarily result from computational and detection delays, as well as inefficiencies associated with the sequential execution of tasks, as discussed in detail later.
@@
To evaluate the potential for performance and improvement in more detail, Figure 27 provides a statistical time analysis based on the subtasks of autonomous operation. The missions discussed in Section 5 are ApproachPallet, LoadPallet, ApproachSlot, and UnloadPallet. The two submissions FindPallets and SelectPallets are included in ApproachPallet to increase readability. The two Approach missions consist mostly of typical navigation movement, whereas the load and unload missions deal mostly with manipulation.
@@
Figure 27: Time evaluation of the sub-missions for ground-to-ground (G to G), ground-to-truck (G to T), and truck-to-ground (T to G) loading for autonomous operation (red) and the forklift operator (cyan).
@@
For the ApproachPallet, a longer time and a higher variance for ADAPT can be observed. This arises mostly from two properties of the pallet detection: Firstly, when no pallets are known to the system, ADAPT stops for several seconds in front of the loading zone to obtain a stable object list. Secondly, as the position and orientation of the pallets are not known in advance, a new approach of pallets is sometimes mandatory for correct loading which takes up to 30 seconds. This occurred multiple times, especially for the ground-to-ground scenario, resulting in a larger variance for the approach. In the ground-to-truck case, the performance of ADAPT and the expert is very similar because the expert operator chooses a suboptimal path, as detailed in the latter. For the ApproachSlot mission, ADAPT achieves a very similar performance, even though more changes in driving direction were made.
@@
For the manipulation subtasks, ADAPT performs very consistently for all scenarios. For loading, the performance of ground-to-ground and ground-to-truck is very similar to that of the human operator. The loading of the truck takes longer for ADAPT. This mostly arises from the time needed to raise the forks to the height of the loading platform and the additional pallet detection check when the pallet is located in a plane different from the movement plane. The unloading of the pallets is close to a standardized task and thus can be done with accurate timing with hardly any variance by the autonomous system. The key difference from manual operation is that the autonomous system raises the forks while stationary, as shown in Figure 21, while a human operator simultaneously manipulates the forks and drives to maximize efficiency. This illustrates a case where ADAPT is not yet optimized for timing performance.
@@
Manual interventions: During 233 minutes of autonomous operation, 60 pallets were loaded and only 12 minutes were spent in manual control, accounting for less than 5% of the total operation time. Most of this manual intervention, 10 minutes, was dedicated to resolving a GNSS localization issue, which is primarily a prototype-related challenge. All other combined manual interventions accounted for just 2 minutes, representing less than 1% of the total operation time. Fig. 30 presents the time analysis of manual interventions during autonomous operation.
@@
To analyze the manual interventions in more detail, three categories of expertise for resolving are introduced: 1) Worker for a typical worker on-site with little knowledge of the machine; 2) Operator, denoting an educated forklift operator capable of driving the machine manually; and 3) Engineer, as a robotic engineer with in-depth knowledge of the autonomy framework. The right-hand side of Fig. 30 shows that most of the time was needed for the Engineer level, i.e., the above mentioned 10 minutes.
In absolute numbers, a total of 17 manual interventions were required. These manual interventions were classified into Collision Detection, Manipulation, and Navigation. The distribution of types of manual intervention is illustrated on the left-hand side of Figure 31.
Figure 31: Overall manual interventions for the autonomous loading operation categorized into type and severity.
Notably, nearly 60% of these interventions were attributed to the collision avoidance module, which requires human clearance for any detected potential collision. Many of these collision warnings occurred near the boundary of the map, where the distance from the collision is shorter if approaching an obstacle or boundary of the map. This issue, resulting from stringent safety constraints in the laboratory environment, can be alleviated by implementing obstacle-classification-based warning distances, assigning larger safety margins for humans and moving obstacles, and smaller margins for static obstacles. Additionally, collision warnings were correctly triggered when humans, including the safety operator, were detected in the vehicle’s path, although such incidents were rare. The second category involves inaccuracies in the detection and manipulation modules, which result in failed pallet load and unload attempts. Implementing recovery routines could reduce the number of necessary manual interventions, albeit at the cost of longer execution times. Lastly, navigation and localization interventions occur when no path to a desired position can be planned, or the localization module fails. On the right-hand side of Figure 31, the severity of manual interventions is classified. Contrary to the time requirement, most human interventions can be solved by a regular worker, commonly available at construction sites. In this case, simply ensuring that the corridor in front of the autonomous vehicle is free of obstacles is sufficient. For ADAPT, this is verified by pressing a confirmation button once. Approximately one-third of interventions require an operator for forklift operations to assist in navigation or manipulation. In our tests, there was one instance where an engineer with knowledge of autonomous functions was needed, due to a failure in the GNSS localization system. These results highlight the substantial reduction in manual labor required for material transport, demonstrating the efficiency of ADAPT.
@@ 7 Conclusion and Outlook
This article presented the autonomous outdoor forklift ADAPT, a system capable of operating in the complex and unstructured environments of construction sites, even in challenging conditions, including low to medium rain. By integrating a forklift-specific sensor suite with a tailored software stack, the system achieves fully autonomous pallet loading and transportation. Extensive real-world testing demonstrated that the autonomous forklift operates at near-human efficiency, achieving more than 80% of expert-level performance in the demonstrated scenarios, while requiring minimal operator intervention.
A key contribution of this work is the novel factor-graph-based joint optimization approach for vehicle localization and pallet mapping, specifically tailored for object manipulation tasks such as pallet loading. This method enhances the system’s adaptability and accuracy in dynamic construction site environments. Additionally integrating a novel fork contact measurement based on pressure feedback significantly improves robustness and safety during object manipulation while remaining cost-effective and simple to implement.
A detailed evaluation compared the autonomous forklift’s performance to that of an expert human operator with over 20 years of experience. The analysis provided insights into performance, robustness, and the nature and frequency of required human interventions. These findings highlight the system’s potential to improve efficiency and safety in material handling for construction sites, addressing labor shortages and reducing operational risks.
Future improvements will focus on motion planning in dynamic environments to enhance safety while increasing operational efficiency by reducing the number of manual interventions. Additionally, the incorporation of semantic information into the environment mapping approach will complement the existing geometric model, allowing the forklift to prioritize stable surfaces such as concrete roads over more challenging terrains like gravel or soil. Further efforts will be directed toward increasing robustness when navigating through challenging terrain and extreme weather conditions. Recurring hardware updates, particularly those involving processing components, will allow replacing existing hardware with more compact, resource-efficient alternatives or utilizing higher-performance components to improve overall timing-related performance. By addressing these aspects, the system can continue to evolve toward a fully integrated, intelligent solution for autonomous material handling in demanding construction environments.